OpenStudy (princeevee):

Find the number halfway between 1/2 and 5/6 A. 27/40 B. 2/3 C. 7/10 D. 13/20 (((will medal plz)))

OpenStudy (princeevee):

anyone?

OpenStudy (anonymous):

it is 2/3 because: 5/6=0.833333333 1/5=0.5 0.8333333333-0.5=0.3333333333 0.333333333/2=0.1666666666666 0.5+0.16666666666=.666666666666 What I did was found what 5/6 and 1/2 was as a decimal because, for me at least and in this case it is easier to work with decimals. So After I got that, I subtracted the smaller # from the bigger # to get the difference. I did this to find how big the gab is between 5/6 and 1/2. After that I devided the gab or .33333333 by 2 to find what the midpoint was and added it to the smaller number to get the midpoint. So your answer is 0.66666666666666 or 2/3

OpenStudy (anonymous):

Hope this helped!

OpenStudy (princeevee):

thx

OpenStudy (princeevee):

can you help me with more?

OpenStudy (anonymous):

How many? Am I doing your homework?

OpenStudy (princeevee):

no

OpenStudy (princeevee):

i just need hlp

OpenStudy (anonymous):

okay.